### v1

Abstract
This paper investigates the continuous-time limit of score-driven models with long memory. By extending score-driven models to incorporate infinite-lag structures with coefficients exhibiting heavy-tailed decay, we establish their weak convergence, under appropriate scaling, to fractional Ornstein-Uhlenbeck processes with Hurst parameter $H<1/2$. When score-driven models are used to characterize the dynamics of volatility, they serve as discrete-time approximations for rough volatility. We present several examples, including EGARCH($\infty$) whose limits give rise to a new class of rough volatility models. Building on this framework, we carry out numerical simulations and option pricing analyses, offering new tools for rough volatility modeling and simulation.
Keywords: Score-driven models, Rough volatility, Weak convergence, Option pricing
2020 MSC: 60F17, 60G22, 62M10, 91B70
1 Introduction
Score-driven models, also known as generalized autoregressive score models, were proposed by Creal et al., (2013) and Harvey, (2013) as a class of observation-driven time series models in which parameter updates are driven by the score—the gradient of the log-likelihood function. By directly linking the score of the data distribution to parameter dynamics in a unified framework, these models inherit flexibility and generality, encompassing many classical models such as GARCH (Bollerslev, , 1986) and ACD (Engle and Russell, , 1998) models. Theoretically, score-driven models have been proven optimal in minimizing the Kullback-Leibler divergence between the true distribution and the postulated distribution (Gorgi et al., , 2024). Moreover, in statistical inference, the model’s closed-form likelihood function facilitates efficient maximum likelihood estimation, exhibiting lower computational costs compared to parameter-driven models. In application, they also demonstrate robust empirical performance across diverse fields such as economics, finance and biology. Within just a few years, the literature on SD models has grown to nearly 400 and can be found online at http://www.gasmodel.com/gaspapers.htm.
As a remarkably influential model in time series analysis, it naturally raises the question: what would its continuous-time counterpart look like? Buccheri et al., (2021) have investigated its continuous-time limit, obtaining a two-dimensional diffusion process. Similarly, Wu and He, (2024) investigated the continuous-time limit of the quasi score-driven volatility model proposed by Blasques et al., (2023), and derived a correlated stochastic volatility model. However, the score-driven models they studied only incorporate first-order or finite-order lags—these are all Markov processes. In fact, when the model was first introduced in Creal et al., (2013), the question was raised as to whether it could be extended to an infinite-lag form. Janus et al., (2014) and Opschoor and Lucas, (2019) pursued such extensions by proposing long memory score-driven models. This naturally leads to the question: What would be the continuous-time counterpart of such a long memory version?
On the other hand, since Gatheral et al., (2018) discovered that volatility is rough, rough volatility models have gained popularity in financial modeling. These models are characterized by the use of fractional Brownian motion with a Hurst parameter $H<1/2$, whose paths are rougher than those of Brownian motion—hence the name. El Euch et al., (2018) investigated the microstructure of rough volatility by employing Hawkes processes to model the arrival dynamics in limit order books. This theoretical basis originates from the discovery of Jaisson and Rosenbaum, (2016) that the scaling limit of nearly unstable heavy-tailed Hawkes processes is a rough Cox-Ingersoll-Ross (CIR) process. It is worth noting that Hawkes processes are self-exciting (similar to observation-driven mechanisms) and exhibit long memory, which provides inspiration for us. In particular, Cai et al., (2024) studied the discrete-time counterpart of the Hawkes process—the INAR($\infty$) process—and obtained similar convergence results. This motivates our investigation into the scaling limit of score-driven models when extended to incorporate long memory.
This paper establishes that a sequence of long memory score-driven models, under appropriate scaling, converges weakly to a rough Ornstein-Uhlenbeck (OU) process. By modeling time-varying parameters as log-volatility (as in EGARCH-type models), we obtain the limit a class of rough volatility models. Consequently, long memory score-driven volatility models can serve as discrete-time approximations for such rough volatility models, facilitating efficient Monte Carlo simulations for financial applications. To the best of our knowledge, apart from Wang and Cui, (2025)’s approach using INAR($\infty$) processes to approximate the rough Heston model, this study presents one of the few methodologies employing time series models to approximate rough volatility processes. Moreover, we present specific examples of long memory score-driven volatility models to further demonstrate the practical versatility and broad applicability of our theoretical results.
The remainder of the paper is structured as follows. Section 2 introduces the long memory score-driven models that we focus on and provides a heuristic derivation of their continuous-time limits. In Section 3, we establish the main convergence results and presents their proofs. Section 4 applies these results to volatility modeling, illustrating with the Gamma-GED-EGARCH($\infty$) model and its extension. We conducts corresponding approximations of rough volatility models and option pricing. Section 5 concludes our findings and provides an outlook for future research.
2 Long memory score-driven models
Since Engle, (1982) discovered that economic data exhibit conditional heteroskedasticity, it has gradually been recognized that certain characteristics of time series are not immutable. Specifically, let $\{y_{n}\}_{n\in\mathbb{N}}$ be a time series and $\mathcal{F}{n}=\sigma(y_{n},y_{n-1},\ldots,y_{0})$ be the $\sigma$-algebra it generates. The conditional density
$y_{n}|\mathcal{F}_{n-1}\sim p(y_{n}|\lambda_{n},\mathcal{F}_{n-1}),$
with a time-varying parameter $\lambda_{n}$. To address this issue, Cox et al., (1981) proposed two modeling approaches: observation-driven models, where the dynamics of the time-varying parameters are governed by past observations of the series itself, and parameter-driven models, in which the parameter dynamics are driven by another stochastic process.
The score-driven model is a class of observation-driven models whose key idea is to update time-varying parameters based on the score of the observation. Specifically, the parameter update follows the form:
$\lambda_{n+1}=\omega+\beta\lambda_{n}+\alpha S(\lambda_{n})\nabla_{n}.$ (2.1)
where $\nabla_{n}=\frac{\partial\log p(y_{n}|\lambda_{n},\mathcal{F}{n-1})}{\partial\lambda_{n}}$ denotes the partial derivative of the log-density function of observations with respect to the parameter, commonly referred to as the (Fisher) score in statistics. Consequently, these models are aptly termed score-driven models. The function $S(\cdot)$ is known as the scaling function, which accounts for the curvature of the log-density function. It is typically chosen as the negative exponent of the conditional Fisher information, i.e., $S(\lambda_{n})=[\mathbb{E}(\nabla_{n}^{2}|\mathcal{F}{n-1})]^{-a}$, with common choices for $a$ being 0, $1/2$, or $1$. To illustrate, when the density function $p(\cdot)$ is chosen to be that of a normal distribution and the parameter $\lambda$ is identified as the conditional variance $\sigma^{2}$, with $a=1$, the score-driven model reduces to the classical GARCH model.
The equation given by (2.1) represents a first-order score-driven model, as the relationship between the score and the parameter is lagged by only one period. This framework can be extended to higher orders, such as $p$-th order or even infinite order, as in Janus et al., (2014),
$\lambda_{n}=\omega+\sum_{i=1}^{n}\phi_{i}S(\lambda_{n-i})\nabla_{n-i}.$
Noting that our time index starts from 0, the summation at time $n$ includes only $n$ terms. The term “infinite” refers to the dependence on the entire history. Additionally, we omit the autoregressive term, as it can be absorbed into the coefficients $\phi_{i}$ through the moving average representation. This formulation appears similar to the ARCH($\infty$) model, but the key difference lies in the fact that the infinite-order lagged term $S\nabla$ in the score-driven model is a martingale difference, whereas the $y^{2}$ term in the ARCH($\infty$) model is not. In other words, rewriting the ARCH($\infty$) model in a form driven by martingale differences gives
$\sigma^{2}_{n}=\omega+\sum_{i=1}^{n}\phi_{i}y^{2}_{n-i}=\omega+\sum_{i=1}^{n}\phi_{i}\sigma^{2}_{n-i}+\sum_{i=1}^{n}\phi_{i}(y^{2}_{n-i}-\sigma^{2}_{n-i}),$
which includes an additional infinite autoregressive component. To address this, Opschoor and Lucas, (2019) proposed the following formulation:
$\lambda_{n}=\omega+\sum_{i=1}^{n}\phi_{i}[\lambda_{n-i}+S(\lambda_{n-i})\nabla_{n-i}].$ (2.2)
The long memory score-driven model we considered henceforth refers to (2.2). To investigate its scaling limit, we define a sequence of processes:
$\begin{split}&y_{t}^{(n)}|\mathcal{F}_{t-1}^{(n)}\sim p\left(y_{t}^{(n)}|\lambda_{t}^{(n)},\mathcal{F}^{(n)}_{t-1}\right),\\ &\lambda_{t}^{(n)}=\omega_{n}+\sum_{i=1}^{t}\phi_{i}^{(n)}\left[\lambda_{t-i}^{(n)}+S(\lambda_{t-i}^{(n)})\nabla_{t-i}^{(n)}\right],\end{split}$ (2.3)
where $\phi_{i}^{(n)}=a_{n}\phi_{i}>0$. Let $a_{n}>0$ and $\uparrow 1$ as $n\to\infty$, $\|\phi\|=\sum_{i=1}^{\infty}\phi_{i}=1$, so that $\|\phi^{(n)}\|\uparrow 1$. This case is referred to as nearly unstable. Specifically, we set $\phi_{i}\sim K/i^{1+\alpha}$, exhibiting power-law decay, where $\alpha\in(1/2,1)$, a form referred to as “heavy-tailed” in Jaisson and Rosenbaum, (2016). This condition is critical for the limit to generate roughness. The sigma-algebra is defined as $\mathcal{F}^{(n)}_{t}=\sigma(\lambda_{0}^{(n)},y_{i}^{(n)},\lambda_{i+1}^{(n)},i=0,1,\ldots,t)$, since $\lambda_{t+1}^{(n)}$ is determined when $\{y_{i}^{(n)}\}_{i\leq t}$ are known.
Our primary focus is on the dynamics of time-varying parameter $\lambda_{t}^{(n)}$, and we aim to study its scaling limiting behavior, which we expect to converge to a SDE driven by a rough fractional Brownian motion. To build intuition for this convergence, we first rewrite the equation. For notation simplicity, the superscript $n$ is sometimes omitted without ambiguity, and the terms $a_{n}$, $\omega_{n}$ serving as a reminder that we are dealing with a sequence indexed by $n$. Define $M_{t}=\sum_{i=0}^{t}S(\lambda_{i})\nabla_{i}$, with $M_{-1}=0$. Clearly, $M$ is a martingale, since
$\mathbb{E}(M_{t}-M_{t-1}|\mathcal{F}_{t-1})=\mathbb{E}[S(\lambda_{t})\nabla_{t}|\mathcal{F}_{t-1}]=S(\lambda_{t})\mathbb{E}[\nabla_{t}|\mathcal{F}_{t-1}]=0.$
Then,
$\begin{split}\lambda_{t}&=\omega_{n}+\sum_{i=0}^{t-1}a_{n}\phi_{t-i}[\lambda_{i}+S(\lambda_{i})\nabla_{i}]\\ &=\omega_{n}+\sum_{i=0}^{t-1}a_{n}\phi_{t-i}(M_{i}-M_{i-1})+\sum_{i=0}^{t-1}a_{n}\phi_{t-i}\lambda_{i}\\ &=\omega_{n}+\sum_{i=0}^{t-1}a_{n}\phi_{t-i}(M_{i}-M_{i-1})+\sum_{i=0}^{t-1}\psi_{t-i}^{(n)}\left[\omega_{n}+\sum_{j=0}^{i-1}a_{n}\phi_{i-j}(M_{j}-M_{j-1})\right].\end{split}$
The final equality follows from the transformation between AR($\infty$) and MA($\infty$), that is,
$\lambda_{t}=\varepsilon_{t}+\sum_{i=0}^{t-1}a_{n}\phi_{t-i}\lambda_{i}\quad\Longleftrightarrow\quad\lambda_{t}=\varepsilon_{t}+\sum_{i=0}^{t-1}\psi^{(n)}_{t-i}\varepsilon_{i}.$
Here, $\psi^{(n)}=\sum_{k=1}^{\infty}(a_{n}\phi)^{*k}$ represents the sum of the $k$-fold convolution of $\phi^{(n)}$. By rearranging the expression and exchanging the order of the double summation, we obtain
$\lambda_{t}=\omega_{n}+\sum_{i=0}^{t-1}\psi_{t-i}^{(n)}\omega_{n}+\sum_{i=0}^{t-1}a_{n}\phi_{t-i}(M_{i}-M_{i-1})+\sum_{j=0}^{t-2}(M_{j}-M_{j-1})\sum_{i=j+1}^{t-1}a_{n}\phi_{i-j}\psi_{t-i}^{(n)}.$
Note that
$\sum_{i=j+1}^{t-1}a_{n}\phi_{i-j}\psi_{t-i}^{(n)}=(\phi^{(n)}*\psi^{(n)})_{t-j},$
and
$\phi^{(n)}*\psi^{(n)}=\sum_{k=1}^{\infty}(\phi^{(n)})^{*{k+1}}=\sum_{k=2}^{\infty}(\phi^{(n)})^{*{k}}=\psi^{(n)}-\phi^{(n)},$
we have
$\begin{split}\lambda_{t}&=\omega_{n}+\sum_{i=0}^{t-1}\psi_{t-i}^{(n)}\omega_{n}+a_{n}\phi_{1}(M_{t-1}-M_{t-2})+\sum_{j=0}^{t-2}\psi_{t-j}^{(n)}(M_{j}-M_{j-1})\\ &=\omega_{n}+\sum_{i=0}^{t-1}\psi_{t-i}^{(n)}\omega_{n}+\sum_{j=0}^{t-1}\psi_{t-j}^{(n)}(M_{j}-M_{j-1}).\end{split}$ (2.4)
Since $\mathbb{E}[\lambda_{t}^{(n)}]\to\omega_{n}/(1-a_{n})$ as $t\to\infty$, we apply a rescaling to $\lambda_{t}^{(n)}$ and define a new process
$\Lambda_{t}^{(n)}:=\frac{(1-a_{n})\theta}{\omega_{n}}\lambda_{\lfloor nt\rfloor}^{(n)},\quad t\in[0,1],$
where $\theta>0$ is a constant controlling the mean level of $\Lambda_{t}^{(n)}$. We aim to investigate the limiting behavior of $\Lambda_{t}^{(n)}$ as $n\to\infty$.
Let $\Delta M_{k}=M_{k}-M_{k-1}$. Assuming
$\mbox{Var}[\Delta M_{k}|\mathcal{F}_{k-1}]=S^{2}(\lambda_{k})\mathbb{E}[(\nabla_{k})^{2}|\mathcal{F}_{k-1}]=U^{2}(\lambda_{k}).$
It follows from (2.4) that
$\begin{split}\Lambda_{t}^{(n)}&=(1-a_{n})\theta+(1-a_{n})\theta\sum_{i=0}^{\lfloor nt\rfloor-1}\psi_{\lfloor nt\rfloor-i}^{(n)}+\frac{1-a_{n}}{\omega_{n}}\theta\sum_{j=0}^{\lfloor nt\rfloor-1}\psi_{\lfloor nt\rfloor-j}^{(n)}\Delta M_{j}^{(n)}\\ &=(1-a_{n})\theta+\theta\sum_{i=1}^{\lfloor nt\rfloor}(1-a_{n})\psi_{i}^{(n)}+\frac{\theta}{\sqrt{n}\omega_{n}}\sum_{j=0}^{\lfloor nt\rfloor-1}n(1-a_{n})\psi_{\lfloor nt\rfloor-j}^{(n)}U(\lambda_{j}^{(n)})\frac{\Delta M_{j}^{(n)}}{U(\lambda_{j}^{(n)})\sqrt{n}}.\end{split}$ (2.5)
Conceivably, when $n\to\infty$,
$\sum_{j=0}^{\lfloor nt\rfloor-1}\frac{\Delta M_{j}^{(n)}}{U(\lambda_{j}^{(n)})\sqrt{n}}\Rightarrow W_{t},$
where $\{W_{t}\}_{t\geq 0}$ is a standard Brownian motion. And according to Cai et al., (2024), in that case $1/2<\alpha<1$, when $1-a_{n}\sim n^{-\alpha}$ tends to zero and the coefficients $\phi_{i}\sim K/i^{1+\alpha}$ exhibit power-law decay, the following weak convergence holds:
$n(1-a_{n})\psi^{(n)}_{\lfloor nx\rfloor}\to f^{\alpha,\kappa}(x):=\kappa x^{\alpha-1}E_{\alpha,\alpha}(-\kappa x^{\alpha}),\quad\mbox{as}\ n\to\infty,$ (2.6)
where $\kappa>0$ is a constant, and
$E_{\alpha,\beta}(x)=\sum_{n=0}^{\infty}\frac{x^{n}}{\Gamma(\alpha n+\beta)}$
is the Mittag-Leffler function. A property of the function is that it decays to zero as $x\to$-$\infty$, and $E_{\alpha,\alpha}(0)=\frac{1}{\Gamma(\alpha)}$. Thus, in terms of singularity, the behavior of $f^{\alpha,\kappa}(t-s)$ is analogous to that of the Riemann-Liouville kernel $(t-s)^{\alpha-1}$.
Therefore, informally and intuitively, as $n\to\infty$, the limit of (2.5) satisfies the following SDE:
$\Lambda_{t}=\theta\int_{0}^{t}f^{\alpha,\kappa}(s){\rm d}s+\nu\int_{0}^{t}f^{\alpha,\kappa}(t-s)U(\Lambda_{s}){\rm d}W_{s}.$ (2.7)
In fact, similar to Proposition 4.10 in El Euch et al., (2018), the SDE (2.7) has the same solution as the SDE
$\Lambda_{t}=\frac{1}{\Gamma(\alpha)}\int_{0}^{t}(t-s)^{\alpha-1}\kappa(\theta-\Lambda_{s}){\rm d}s+\frac{\kappa\nu}{\Gamma(\alpha)}\int_{0}^{t}(t-s)^{\alpha-1}U(\Lambda_{s}){\rm d}W_{s}.$ (2.8)
Since $\alpha-1\in(-1/2,0)$, the term $\frac{1}{\Gamma(\alpha)}\int_{0}^{t}(t-s)^{\alpha-1}{\rm d}W_{s}$ corresponds to a fractional Brownian motion with Hurst parameter $H<1/2$, which means (2.8) falls into the rough case. It is worth mentioning that when $U(x)=\sqrt{x}$, this actually corresponds to the rough fractional CIR process described in Jaisson and Rosenbaum, (2016).
3 Main results
Following the heuristic derivation above, in this section, we state our main theorem and present its proof. We begin with the following assumptions.
Assumption 1.
There exists a constant $\gamma<\infty$, such that $U(x)\to\gamma$, as $x\to\infty$.
Assumption 2.
The sequences $a_{n}$ and $\omega_{n}$ have the following asymptotic behavior as $n\to\infty$:
$1-a_{n}\sim n^{-\alpha},\quad\omega_{n}\sim n^{-\frac{1}{2}}.$
Assumption 3.
For each $n$, the martingale difference sequence $\{\Delta M_{k}^{(n)}\}_{k\in\mathbb{N}}$ is strictly stationary. Furthermore, for any $p\geq 2$,
$\sup_{n\in\mathbb{N}}\mathbb{E}\left|\Delta M_{k}^{(n)}\right|^{p}<\infty.$
Then, we have the following theorem.
Theorem 1.
Under Assumptions 1-3, the rescaled long memory score-driven parameter process
$\Lambda_{t}^{(n)}:=\frac{1-a_{n}}{\omega_{n}}\theta\lambda^{(n)}_{\lfloor nt\rfloor},\quad t\in[0,1]$
converges weakly in the Skorohod space $\mathcal{D}([0,1])$ to a rough fractional diffusion process given by the following SDE:
$\Lambda_{t}=\frac{1}{\Gamma(\alpha)}\int_{0}^{t}(t-s)^{\alpha-1}\kappa(\theta-\Lambda_{s}){\rm d}s+\frac{\gamma\kappa\nu}{\Gamma(\alpha)}\int_{0}^{t}(t-s)^{\alpha-1}{\rm d}W_{s}.$ (3.1)
Where $\kappa=\lim_{n\to\infty}\frac{\alpha n^{\alpha}(1-a_{n})}{K\Gamma(1-\alpha)}$ and $\nu=\lim_{n\to\infty}\frac{\theta}{\sqrt{n}\omega_{n}}$. The existence of these limits is guaranteed by Assumption 2.
In fact, we restrict our attention to the case of rough OU process in (2.8). This is because our proof strategy fundamentally differs from that of Jaisson and Rosenbaum, (2016) or Wang and Cui, (2025): while they show that the Hawkes process (equivalent to the integrated intensity in probability) converges to a integrated rough CIR process, we directly establish the convergence of the parameter process itself to a rough process, rather than its integral. The difficulty in dealing with the process itself lies in verifying its tightness and we attempt to circumvent this by applying the continuous mapping theorem.
To this end, we regard the rough diffusion process as the image of a diffusion process under the generalized fractional operator (GFO), which is inspired by Horvath et al., (2024). Specifically, for $\lambda\in(0,1)$ and $\alpha\in(-\lambda,1-\lambda)$, $\mathcal{G}^{\alpha}$ is the GFO associated to kernel
$g\in\mathcal{L}^{\alpha}:=\left\{g\in C^{2}((0,1]):\left|\frac{g(u)}{u^{\alpha}}\right|,\left|\frac{g^{\prime}(u)}{u^{\alpha-1}}\right|\ \text{and}\ \left|\frac{g^{\prime\prime}(u)}{u^{\alpha-2}}\right|\ \mbox{are bounded}\right\}$
defined on $\lambda$-Hölder space $\mathcal{C}^{\lambda}([0,1])$,
$(\mathcal{G}^{\alpha}f)(t)=\left\{\begin{aligned} &\int_{0}^{t}(f(s)-f(0))\frac{{\rm d}}{{\rm d}t}g(t-s){\rm d}s,\quad\text{if}\ \alpha\geq 0,\\ &\frac{{\rm d}}{{\rm d}t}\int_{0}^{t}(f(s)-f(0))g(t-s){\rm d}s,\quad\text{if}\ \alpha<0.\end{aligned}\right$.
When $g(x)=x^{\alpha}$, this reduces to the classical Riemann-Liouville operator. In this paper, we choose $g(x)=f^{\alpha,\kappa}(x)\in\mathcal{L}^{\alpha-1}$ and focus on the case $\alpha\in(1/2,1)$. Consequently, the corresponding GFO actually becomes
$(\mathcal{G}^{\alpha-1}f)(t)=\frac{{\rm d}}{{\rm d}t}\int_{0}^{t}(f(s)-f(0))f^{\alpha,\kappa}(t-s){\rm d}s.$ (3.2)
Define $Y_{t}:=\theta t+\nu\gamma W_{t}$, and
$\Lambda_{t}=\theta\int_{0}^{t}f^{\alpha,\kappa}(s){\rm d}s+\nu\gamma\int_{0}^{t}f^{\alpha,\kappa}(t-s){\rm d}W_{s}$
is actually has the same solution as the (3.1). Consider the partition $\mathcal{T}:=\{t_{i}=i/n,i=0,1,\ldots,n\}$. We then define
$Y_{t}^{(n)}=\theta t+\frac{\theta}{n\omega_{n}}\left[(1-nt+\lfloor nt\rfloor)M_{\lfloor nt\rfloor-1}^{(n)}+(nt-\lfloor nt\rfloor)M_{\lfloor nt\rfloor}^{(n)}\right],$ (3.3)
$\widetilde{\Lambda}_{t}^{(n)}=\sum_{i=0}^{\lfloor nt\rfloor-1}\left(\theta+\frac{\theta}{\omega_{n}}\Delta M_{i}^{(n)}\right)\int_{t_{i}}^{t_{i+1}}f^{\alpha,\kappa}(t-s){\rm d}s+\left(\theta+\frac{\theta}{\omega_{n}}\Delta M_{\lfloor nt\rfloor}^{(n)}\right)\int_{t_{\lfloor nt\rfloor}}^{t}f^{\alpha,\kappa}(t-s){\rm d}s.$ (3.4)
The subsequent proof will proceed in four steps, as illustrated by the following diagram, which provides a clear path.
  1. (i)
    Establishing the representation $\Lambda_{t}=\mathcal{G}^{\alpha-1}Y_{t}$, and $\widetilde{\Lambda}_{t}^{(n)}=\mathcal{G}^{\alpha-1}Y_{t}^{(n)}$;
  2. (ii)
    Proving the weak convergence $Y^{(n)}\Rightarrow Y$ in Hölder space $\mathcal{C}^{\lambda}$ for $\lambda<1/2$;
  3. (iii)
    Demonstrating that $\mathcal{G}^{\alpha-1}$ is a continuous operator from $\mathcal{C}^{\lambda}$ to $\mathcal{C}^{\lambda+\alpha-1}$, which yields the convergence $\widetilde{\Lambda}_{t}^{(n)}\Rightarrow\Lambda_{t}$ in $\mathcal{C}^{\lambda+\alpha-1}$, thus in $\mathcal{D}$;
  4. (iv)
    Finally, by proving $\widetilde{\Lambda}_{t}^{(n)}$ and $\Lambda_{t}^{(n)}$ are asymptotically indistinguishable, we obtain $\Lambda_{t}^{(n)}\Rightarrow\Lambda_{t}$.
$Y^{(n)}_{t}$$Y_{t}$$(\mathcal{C}^{\lambda}([0,1]),\|\cdot\|_{\lambda})$$\Lambda^{(n)}_{t}$$\widetilde{\Lambda}^{(n)}_{t}$$\Lambda_{t}$$(\mathcal{C}^{\lambda+\alpha-1}([0,1]),\|\cdot\|_{\lambda+\alpha-1})$$\scriptstyle{\|\cdot\|_{\lambda}}$$\scriptstyle{\mathcal{G}^{\alpha-1}}$$\scriptstyle{\mathcal{G}^{\alpha-1}}$$\scriptstyle{\mathcal{G}^{\alpha-1}}\scriptstyle{\approx}$$\scriptstyle{\|\cdot\|_{\lambda+\alpha-1}}$
Before proceeding with the main proof, we first recall the key convergence result (2.6) for the Mittag-Leffler function, and extend this result to $L^{2}$-convergence for our subsequent arguments.
3.1 Converges to $f^{\alpha,\kappa}(x)$
Jaisson and Rosenbaum, (2016) prove that the mean of geometric sums converges in distribution to a random variable with density function $f^{\alpha,\kappa}(x)$. Our setting differs slightly because the summed terms follow a discrete distribution.
Specifically, we define the sequence of random variables
$G_{n}=\frac{\sum_{k=1}^{I_{n}}U_{k}}{n},$
where $I_{n}$ follows a geometric distribution with parameter $1-a_{n}$, that is, $\mathbb{P}(I_{n}=i)=a_{n}^{i-1}(1-a_{n})$; $U_{k}$} is a sequence of i.i.d. random variables with distribution $\phi$, that is, $\mathbb{P}(U_{1}=i)=\phi_{i}$. Then, the random variable $G_{n}$, called the mean of geometric sums, has its law given by
$\begin{split}\mathbb{P}(G_{n}=i)&=\sum_{j=1}^{\infty}\mathbb{P}(\sum_{k=1}^{j}U_{k}=ni)\cdot\mathbb{P}(I_{n}=j)=(1-a_{n})\sum_{j=1}^{\infty}\phi^{*j}_{ni}a_{n}^{j-1}\\ &=\frac{1-a_{n}}{a_{n}}\sum_{j=1}^{\infty}(a_{n}\phi)^{*j}_{ni}=\frac{(1-a_{n})\psi^{(n)}_{ni}}{a_{n}},\quad\text{for}\ ni\in\mathbb{N}^{+}.\end{split}$
Thus, its distribution function is $F_{n}(x)=\sum_{i=1}^{\lfloor nx\rfloor}\frac{(1-a_{n})\psi^{(n)}_{i}}{a_{n}}$, density function $\rho_{n}(x)=\frac{n(1-a_{n})}{a_{n}}\psi^{(n)}_{\lfloor nx\rfloor}$. Considering the characteristic function of $G_{n}$.
$\begin{split}\hat{\rho}_{n}(u)&=\mathbb{E}({\rm e}^{-{\rm i}uG_{n}})=\mathbb{E}({\rm e}^{-\frac{{\rm i}u}{n}\sum_{k=1}^{I_{n}}U_{k}})=\sum_{j=1}^{\infty}\mathbb{E}({\rm e}^{-\frac{{\rm i}u}{n}\sum_{k=1}^{j}U_{k}})\mathbb{P}(I_{n}=j)\\ &=(1-a_{n})\sum_{j=1}^{\infty}[\mathbb{E}({\rm e}^{-\frac{{\rm i}u}{n}U_{k}})]^{j}a_{n}^{j-1}=\frac{1-a_{n}}{a_{n}}\sum_{j=1}^{\infty}[a_{n}\hat{\phi}(u/n)]^{j}\\ &=\frac{1-a_{n}}{a_{n}}\frac{a_{n}\hat{\phi}(u/n)}{1-a_{n}\hat{\phi}(u/n)}=\frac{\hat{\phi}(u/n)}{1-\frac{a_{n}}{1-a_{n}}[\hat{\phi}(u/n)-1]},\end{split}$
where $\hat{\phi}$ denotes the Fourier transform of $\phi$, or equivalently, the characteristic function of the random variable $U_{1}$. Since $\phi_{i}\sim\frac{K}{i^{1+\alpha}}$, by the Karamata-Tauberian theorem (see Cai et al., (2024)), we can characterize the behavior of $\hat{\phi}$ near the origin,
$1-\hat{\phi}(u)\underset{u\to 0}{\sim}\frac{K\Gamma(1-\alpha)}{\alpha}({\rm i}u)^{\alpha}.$ (3.5)
Assumption 2 implies that the limit of $\frac{\alpha n^{\alpha}(1-a_{n})}{K\Gamma(1-\alpha)}$ exists and we denote it by $\kappa$. Consequently, for any fixed $u$, we have
$\hat{\rho}_{n}(u)\to\frac{\kappa}{\kappa+({\rm i}u)^{\alpha}},\quad\text{as}\ n\to\infty,$
and the limit happens to be the Fourier transform of $f^{\alpha,\kappa}(x)=\kappa x^{\alpha-1}E_{\alpha,\alpha}(-\kappa x^{\alpha})$. Therefore,
$F_{n}(x)=\sum_{i=1}^{\lfloor nx\rfloor}\frac{(1-a_{n})\psi^{(n)}_{i}}{a_{n}}\to\int_{0}^{x}f^{\alpha,\kappa}(s){\rm d}s,\quad\text{as}\ n\to\infty.$ (3.6)
By Dini’s theorem, this convergence is uniform.
In fact, for the density function itself, the convergence $\hat{\rho}_{n}\to\hat{f}^{\alpha,\kappa}$ implies only that the density function $\rho_{n}$ converges weakly to $f^{\alpha,\kappa}$. We now proceed to prove that this convergence also holds in $L^{2}$. To this end, we first provide some estimates for $\hat{\phi}$. Note that in $\hat{\rho}_{n}$, the function $\hat{\phi}$ appears as the form $\hat{\phi}(u/n)=:\hat{\phi}_{n}(u)$. We regard this as the Fourier transform of the step function $\phi_{n}(x)=\phi_{\lfloor nx\rfloor}$, which is consistent with the representation of $\rho_{n}$.
Lemma 2.
For any $|u|>1$, we have $|\hat{\phi}_{n}(u)|\leq|u|^{-\alpha}$. Additionally, there exist constants $c_{1},c_{2}>0$ such that
$|1-\hat{\phi}_{n}(u)|\geq\left\{\begin{aligned} c_{1}|u/n|^{\alpha},\quad\text{if}\ |u|\leq 1,\\ c_{2},\quad\text{if}\ |u|>1.\end{aligned}\right$.
Proof.
Since $\lim_{x\to\infty}\phi_{n}(x)=0$, integration by parts yields
$\begin{split}\hat{\phi}_{n}(u)=\int_{0}^{\infty}{\rm e}^{-{\rm i}ux}\phi_{n}(x){\rm d}x=\frac{1}{{\rm i}u}\sum_{k=1}^{\infty}{\rm e}^{-{\rm i}uk}\phi_{nk}.\end{split}$
Since $\|\phi\|=1$, for any $|u|>1$, we have
$|\hat{\phi}_{n}(u)|\leq\frac{1}{|u|}\sum_{k=1}^{\infty}\phi_{nk}\leq\frac{1}{|u|}\leq\frac{1}{|u|^{\alpha}}.$
From (3.5), there exists $\delta<1$ such that for all $|u|<\delta$, we have
$|1-\hat{\phi}_{n}(u)|\geq\frac{K\Gamma(1-\alpha)}{2\alpha}\left|\frac{u}{n}\right|^{\alpha}.$
Moreover, $\frac{|1-\hat{\phi}_{n}(u)|}{|u/n|^{\alpha}}$ is a continuous function on $\delta,1]$, and thus attains its minimum at some point $u_{0}\in[\delta,1]$. Suppose $|1-\hat{\phi}_{n}(u_{0})|=0$, that is $\text{Re}(\hat{\phi}_{n}(u_{0}))=\mathbb{E}[\cos u_{0}U_{1}/n]=1$. This implies that $U_{1}\in\{2kn\pi/u_{0},k=0,1,2,\ldots\}$ almost surely, but this is clearly impossible. Therefore, there exists $m=\frac{|1-\hat{\phi}(u_{0})|}{|u_{0}/n|^{\alpha}}>0$ such that
$|1-\hat{\phi}_{n}(u)|\geq m\left|\frac{u}{n}\right|^{\alpha}.$
Consequently, for $|u|\leq 1$, taking $c_{1}=\min\left\{\frac{K\Gamma(1-\alpha)}{2\alpha},m\right\}$ yields the desired result.
Similarly, since $|\hat{\phi}_{n}(u)|\leq|u|^{-\alpha}$, there exists $\delta^{\prime}>1$ such that for all $|u|>\delta^{\prime}$, we have $|1-\hat{\phi}_{n}(u)|\geq 1/2$. Moreover, $|1-\hat{\phi}_{n}(u)|$ is continuous on $1,\delta^{\prime}]$, and thus attains its minimum at some point $u_{1}\in[1,\delta^{\prime}]$. Suppose $|1-\hat{\phi}_{n}(u_{1})|=0$, that is $\text{Re}(\hat{\phi}_{n}(u_{1}))=\mathbb{E}[\cos u_{1}U_{1}/n]=1$. This implies that $U_{1}\in\{2kn\pi/u_{1},k=0,1,2,\ldots\}$ almost surely, but this is clearly impossible. Therefore, there exists $m^{\prime}=|1-\hat{\phi}_{n}(u_{1})|>0$ such that
$|1-\hat{\phi}_{n}(u)|\geq m^{\prime}.$
Consequently, for $|u|>1$, taking $c_{2}=\min\left\{1/2,m^{\prime}\right\}$ yields the desired result. ∎
Proposition 3.
The sequence $\rho_{n}$ converges to $f^{\alpha,\kappa}$ in the sense of $L^{2}$, i.e.,
$\|\rho_{n}-f^{\alpha,\kappa}\|_{2}:=\int_{0}^{1}\left[\rho_{n}(x)-f^{\alpha,\kappa}(x)\right]^{2}{\rm d}x\to 0,\ \text{as}\ n\to\infty.$ (3.7)
Proof.
To establish the $L^{2}$-convergence, we employ the Fourier isometry
$\|\rho_{n}-f^{\alpha,\kappa}\|_{2}=\frac{1}{2\pi}\|\hat{\rho}_{n}-\hat{f}^{\alpha,\kappa}\|_{2}.$
which reduces the problem to verifying $L^{2}$-convergence of the corresponding characteristic functions. Given that $\widehat{\rho}_{n}\to\widehat{f}^{\alpha,\kappa}$ pointwise, it suffices to verify that the sequence $\widehat{\rho}_{n}$ satisfies the dominated convergence theorem.
According to estimates in Lemma 2, when $|u|\leq 1$,
$|\hat{\rho}_{n}(u)|=\left|\frac{(1-a_{n})\hat{\phi}_{n}(u)}{1-a_{n}\hat{\phi}_{n}(u)}\right|\leq\frac{|1-a_{n}|}{|1-\hat{\phi}_{n}(u)|}\leq\frac{c_{1}}{|u|^{\alpha}}.$
when $|u|>1$,
$|\hat{\rho}_{n}(u)|=\left|\frac{(1-a_{n})\hat{\phi}_{n}(u)}{1-a_{n}\hat{\phi}_{n}(u)}\right|\leq\frac{|\hat{\phi}_{n}(u)|}{|1-\hat{\phi}(u/n)|}\leq\frac{1}{c_{2}|u|^{\alpha}}.$
Thus, there exists $C=\max\{c_{1},1/c_{2}\}$ such that for all $u\in\mathbb{R}$, we have
$|\hat{\rho}_{n}(u)|\leq 1\wedge\frac{C}{|u|^{\alpha}},$
and the right-hand side
$\int_{\mathbb{R}}\left(1\wedge\frac{C}{|u|^{\alpha}}\right)^{2}{\rm d}u=2\left(\int_{0}^{C^{-\alpha}}1{\rm d}u+\int_{C^{-\alpha}}^{\infty}\frac{C^{2}}{u^{2\alpha}}{\rm d}u\right)=2C^{-\alpha}+\frac{2C^{2\alpha-1}}{1-2\alpha}<\infty$
is $L^{2}$-integrable. The result thus follows by applying the dominated convergence theorem. ∎
3.2 Proof of main theorem
We proceed with the proof following the aforementioned four steps. First, we need to demonstrate how the convergence target $\Lambda_{t}$ can be expressed in terms of the operator $\mathcal{G}^{\alpha-1}$ in (3.2).
Proposition 4.
The equality $\Lambda_{t}=(\mathcal{G}^{\alpha-1}Y)_{t}$ holds almost surely for all $t\in[0,1]$.
Proof.
Since the paths of $Y_{t}$ have $1/2-\epsilon$ Hölder continuity, for $\alpha\in(1/2,1)$, the exponent $\alpha-1$ fall within (1/2,0)(-1/2,0), which is well-defined in this context of GFO. By definition of (3.2),
$(\mathcal{G}^{\alpha-1}Y)(t)=\frac{{\rm d}}{{\rm d}t}\int_{0}^{t}(Y(s)-Y(0))f^{\alpha,\kappa}(t-s){\rm d}s.$
For $\varepsilon>0$, define the operator
$(\mathcal{G}^{\alpha}_{\varepsilon}Y)(t)=\int_{0}^{t-\varepsilon}(Y(s)-Y(0))f^{\alpha,\kappa}(t-s){\rm d}s.$
Then, for any $t\in[0,1]$, we almost surely have
[…]
This convergence is uniform since the $L^{2}$-integrability of $f^{\alpha,\kappa}$. Consequently, we obtain
$(\mathcal{G}^{\alpha-1}Y)(t)=\frac{{\rm d}}{{\rm d}t}\lim_{\varepsilon\to 0}(\mathcal{G}^{\alpha}_{\varepsilon}Y)(t)=\lim_{\varepsilon\to 0}\frac{{\rm d}}{{\rm d}t}(\mathcal{G}^{\alpha}_{\varepsilon}Y)(t)=\Lambda_{t}.$
Proposition 5.
The equality $\widetilde{\Lambda}_{t}^{(n)}=(\mathcal{G}^{\alpha-1}Y^{(n)})_{t}$ holds almost surely for all $t\in[0,1]$.
Proof.
It should be noted that $Y^{(n)}$ is actually a rescaled version of the martingale $M^{(n)}$. To ensure its paths belong to a Hölder space, we adopt the linear interpolation
$Y_{t}^{(n)}=\theta t+\frac{\theta}{n\omega_{n}}\left[(1-nt+\lfloor nt\rfloor)M_{\lfloor nt\rfloor-1}^{(n)}+(nt-\lfloor nt\rfloor)M_{\lfloor nt\rfloor}^{(n)}\right],$
which is piecewise differentiable, and for $t\in(t_{i},t_{i+1})$, it holds that
$\frac{{\rm d}Y^{(n)}(t)}{{\rm d}t}=\theta+\frac{\theta}{\omega_{n}}\Delta M_{i}^{(n)}.$
By the Assumption 3, its paths is Lipschitz continuous, the GFO is well-defined on $Y_{t}^{(n)}$. Thus, by the definition (3.2),
$\begin{split}&\int_{0}^{t}(Y^{(n)}(s)-Y^{(n)}(0))f^{\alpha,\kappa}(t-s){\rm d}s=\int_{0}^{t}\int_{0}^{t-s}f^{\alpha,\kappa}(u){\rm d}u\frac{{\rm d}(Y^{(n)}(s)-Y^{(n)}(0))}{{\rm d}s}{\rm d}s\\ &=\sum_{i=0}^{\lfloor nt\rfloor-1}\left(\theta+\frac{\theta}{\omega_{n}}\Delta M_{i}^{(n)}\right)\int_{t_{i}}^{t_{i+1}}\int_{0}^{t-s}f^{\alpha,\kappa}(u){\rm d}u{\rm d}s+\left(\theta+\frac{\theta}{\omega_{n}}\Delta M_{\lfloor nt\rfloor}^{(n)}\right)\int_{t_{\lfloor nt\rfloor}}^{t}\int_{0}^{t-s}f^{\alpha,\kappa}(u){\rm d}u{\rm d}s.\end{split}$
Thus, by the definition (3.2),
$\begin{split}&(\mathcal{G}^{\alpha-1}Y^{(n)})(t)=\frac{{\rm d}}{{\rm d}t}\int_{0}^{t}(Y^{(n)}(s)-Y^{(n)}(0))f^{\alpha,\kappa}(t-s){\rm d}s\\ &=\sum_{i=0}^{\lfloor nt\rfloor-1}\left(\theta+\frac{\theta}{\omega_{n}}\Delta M_{i}^{(n)}\right)\int_{t_{i}}^{t_{i+1}}f^{\alpha,\kappa}(t-s){\rm d}s+\left(\theta+\frac{\theta}{\omega_{n}}\Delta M_{\lfloor nt\rfloor}^{(n)}\right)\int_{t_{\lfloor nt\rfloor}}^{t}f^{\alpha,\kappa}(t-s){\rm d}s.\end{split}$
The second step, we prove that $Y^{(n)}\Rightarrow Y$ in Hölder space $\mathcal{C}^{\lambda}$ for $\lambda<1/2$. The Donsker invariance principle for convergence in Hölder spaces was first established by Lamperti, (1962). Račkauskas and Suquet, (2004) provided necessary and sufficient conditions for the convergence of i.i.d. partial sum processes in Hölder spaces, while results for martingale difference sequences were derived by Giraudo, (2016) from a dynamical systems perspective.
Proposition 6.
For any $\lambda<1/2$, $Y^{(n)}$ converges weakly to $Y$ in Hölder space $\mathcal{C}^{\lambda}([0,1])$.
Proof.
Note that
$Y_{t}^{(n)}=\theta t+\frac{\theta}{n\omega_{n}}\left(\sum_{i=0}^{\lfloor nt\rfloor-1}\Delta M^{(n)}_{i}+(nt-\lfloor nt\rfloor)\Delta M^{(n)}_{\lfloor nt\rfloor}\right)$,
by Theorem 2.2 in Giraudo, (2016), it suffices to prove that
$\lim_{t\to\infty}t^{p}\mathbb{P}\left(|\Delta M_{k}^{(n)}|>t\right)=0\ \text{and}\ \mathbb{E}\left[|\Delta M_{k}^{(n)}|^{2}|\mathcal{F}_{k-1}\right]\in L^{p/2}.$ (3.8)
Then
$n^{-1/2}\left(\sum_{i=0}^{\lfloor nt\rfloor-1}\Delta M^{(n)}_{i}+(nt-\lfloor nt\rfloor)\Delta M^{(n)}_{\lfloor nt\rfloor}\right)\Rightarrow\sigma W_{t}\ \text{in}\ \mathcal{C}^{1/2-1/p}([0,1]),$
where $\sigma^{2}=\lim_{n\to\infty}\frac{1}{n}\mathbb{E}[\sum_{i=0}^{n-1}\Delta M^{(n)}_{i}]^{2}\in L^{1}$. For $p^{\prime}>0$, if $\mathbb{E}|\Delta M_{k}^{(n)}|^{p^{\prime}}<\infty$, by Markov’s inequality, for any positive $p<p^{\prime}$, we have
$t^{p}\mathbb{P}\left(|\Delta M_{k}^{(n)}|>t\right)\leq t^{p-p^{\prime}}\mathbb{E}\left|\Delta M_{k}^{(n)}\right|^{p^{\prime}}\to 0,\text{as}\ t\to\infty.$
By Jensen’s inequality,
$\mathbb{E}\left(\mathbb{E}\left[|\Delta M_{k}^{(n)}|^{2}|\mathcal{F}_{k-1}\right]\right)^{\frac{p}{2}}\leq\mathbb{E}\left(\mathbb{E}\left[|\Delta M_{k}^{(n)}|^{2\cdot\frac{p}{2}}|\mathcal{F}_{k-1}\right]\right)=\mathbb{E}\left|\Delta M_{k}^{(n)}\right|^{p}<\infty.$
Therefore, Assumption 3 ensures that condition (3.8) holds for any $p>2$. Moreover,
$\lim_{n\to\infty}\frac{1}{n}\mathbb{E}\left[\sum_{i=0}^{n-1}\Delta M^{(n)}_{i}\right]^{2}=\lim_{n\to\infty}\sum_{i=0}^{n-1}\frac{1}{n}\mathbb{E}\left[\mathbb{E}\left(|\Delta M_{i}^{(n)}|^{2}|\mathcal{F}_{i-1}\right)\right]=\lim_{n\to\infty}\sum_{i=0}^{n-1}\frac{1}{n}\mathbb{E}\left[U^{2}(\lambda_{i}^{(n)})\right]=\gamma^{2},$
together with $\frac{\theta}{\sqrt{n}\omega}\to\nu$, we have the weak convergence
$Y_{t}^{(n)}\Rightarrow\theta t+\nu\gamma W_{t}=Y_{t}$
in $\mathcal{C}^{1/2-1/p}([0,1])$, for any $p>2$. ∎
Subsequently, we establish the continuity of the operator $\mathcal{G}^{\alpha-1}$, thereby enabling the application of the continuous mapping theorem. Indeed, this result is encompassed within Proposition 2.2 of Horvath et al., (2024), where a proof of continuity for general GFOs is provided. We directly apply their result to obtain the following proposition.
Proposition 7.
For any $\alpha\in(1/2,1)$, there exists $\lambda\in(1-\alpha,1/2)$ such that the operator $\mathcal{G}^{\alpha}$ is continuous from $\mathcal{C}^{\lambda}([0,1])$ to $\mathcal{C}^{\lambda+\alpha-1}([0,1])$.
This continuity together with the fact $Y^{(n)}\Rightarrow Y$, we can obtain that $\mathcal{G}^{\alpha-1}Y^{(n)}\Rightarrow\mathcal{G}^{\alpha-1}Y$ by continuous mapping theorem. From the above representations, this entails
$\widetilde{\Lambda}_{t}^{(n)}\Rightarrow\Lambda_{t}\ \text{in}\ \mathcal{C}^{\lambda+\alpha-1}([0,1]).$
According to Lemma 3.10 in Horvath et al., (2024), we can readily extend this weak convergence to the continuous function space $\mathcal{C}([0,1])$ and Skorokhod space $\mathcal{D}([0,1])$.
We have established the convergence of $\widetilde{\Lambda}_{t}^{(n)}$. However, our primary process of interest is $\Lambda_{t}^{(n)}$ given by (2.5), with $\widetilde{\Lambda}_{t}^{(n)}$ serving merely as an auxiliary process. Finally, we will demonstrate that the two processes are asymptotically indistinguishable. By virtue of Theorem 3.1 in Billingsley, (2013), it follows that $\Lambda_{t}^{(n)}\Rightarrow\Lambda_{t}$.
Proposition 8.
The two processes $\Lambda_{t}^{(n)}$ and $\widetilde{\Lambda}_{t}^{(n)}$ are asymptotically indistinguishable in $\mathcal{C}([0,1])$, that is,
$\sup_{t\in[0,1]}\left|\Lambda_{t}^{(n)}-\widetilde{\Lambda}_{t}^{(n)}\right|\Rightarrow 0.$
Proof.
By comparing (2.5) with (3.4), we have
$\begin{split}&\left|\Lambda_{t}^{(n)}-\widetilde{\Lambda}_{t}^{(n)}\right|\\ &\leq\left|(1-a_{n})\theta\right|+\theta\left|\sum_{i=1}^{\lfloor nt\rfloor}(1-a_{n})\psi_{i}^{(n)}-\int_{0}^{t}f^{\alpha,\kappa}(u){\rm d}u\right|+\frac{\theta}{\omega_{n}}\left|\Delta M_{\lfloor nt\rfloor}^{(n)}\int_{t_{\lfloor nt\rfloor}}^{t}f^{\alpha,\kappa}(t-s){\rm d}s\right|\\ &+\frac{\theta}{n\omega_{n}}\left|\sum_{i=0}^{\lfloor nt\rfloor-1}\left[n(1-a_{n})\psi_{\lfloor nt\rfloor-i}^{(n)}-n\int_{t_{i}}^{t_{i+1}}f^{\alpha,\kappa}(t-s){\rm d}s\right]\Delta M_{i}^{(n)}\right|=I_{1}+I_{2}+I_{3}+I_{4}.\end{split}$
It is evident that $I_{1}\to 0$. By (3.6) and Dini’s theorem, $I_{2}$ converges to zero uniformly. As for $I_{3}$,
$\mathbb{E}|\sup_{t\in[0,1]}I_{3}|\leq\frac{\theta}{n\omega_{n}}\mathbb{E}\left|\Delta M_{1}^{(n)}\right|\sup_{s\in(0,t-t_{\lfloor nt\rfloor}]}\left|f^{\alpha,\kappa}(s)\right|\leq\frac{C}{n\omega_{n}}\left(\frac{1}{n}\right)^{\alpha-1}\sim n^{1/2-\alpha}\to 0.$
Finally, we focus on $I_{4}$, aiming to prove that its supremum converges to zero in probability. Noting that $n(1-a_{n})\psi_{\lfloor nt\rfloor}^{(n)}/\rho_{n}(t)=a_{n}\to 1$, we will henceforth replace $n(1-a_{n})\psi_{\lfloor nt\rfloor-i}^{(n)}$ with $\rho_{n}(t-t_{i})$. Then, we employ the left-endpoint rule to approximate the integral $\int_{t_{i}}^{t_{i+1}}f^{\alpha,\kappa}(t-s){\rm d}s$, and this approximation is uniformly since $f^{\alpha,\kappa}\in L^{r}$ for $r<1/(1-\alpha)$.
Therefore, for any $\varepsilon>0$,
$\begin{split}\mathbb{P}\left(\sup_{t\in[0,1]}|I_{4}|\geq\varepsilon\right)&\sim\mathbb{P}\left(\sup_{k=1,\ldots,n}\left|\frac{\theta}{n\omega_{n}}\sum_{i=0}^{k-1}\left[\rho_{n}\left(t_{k}-t_{i}\right)-f^{\alpha,\kappa}(t_{k}-t_{i})\right]\Delta M_{i}^{(n)}\right|\geq\varepsilon\right)\\ &\leq\left(\frac{\theta}{\varepsilon n\omega_{n}}\right)^{2}\mathbb{E}\left|\sum_{i=0}^{n-1}\left[\rho_{n}(t_{k}-t_{i})-f^{\alpha,\kappa}(t_{k}-t_{i})\right]\Delta M_{i}^{(n)}\right|^{2}\\ &=\left(\frac{\theta}{\varepsilon n\omega_{n}}\right)^{2}\sum_{i=0}^{n-1}\left[\rho_{n}(t_{k}-t_{i})-f^{\alpha,\kappa}(t_{k}-t_{i})\right]^{2}\mathbb{E}\left[\Delta M_{i}^{(n)}\right]^{2}\\ &=\frac{\theta^{2}\sup_{n}\mathbb{E}\left[\Delta M_{i}^{(n)}\right]^{2}}{\varepsilon^{2}n\omega_{n}^{2}}\sum_{i=1}^{n}\left[\rho_{n}(t_{i})-f^{\alpha,\kappa}(t_{i})\right]^{2}\frac{1}{n}\\ &\lesssim C_{\varepsilon}\|\rho_{n}-f^{\alpha,\kappa}\|_{2}^{2}\end{split}$
By Proposition 3, it follows that the above probability converges to zero. Consequently, $\Lambda_{t}^{(n)}$ and $\widetilde{\Lambda}_{t}^{(n)}$ are asymptotically indistinguishable. ∎
4 Specific examples for volatility models
4.1 Gamma-GED-EGARCH($\infty$) model
When we use the score-driven model to characterize volatility, we focus on the following observed time series:
$y_{n}=\sqrt{\varphi(\lambda_{n})}\varepsilon_{n},\quad\varepsilon_{n}|\mathcal{F}_{n-1}\stackrel{{\scriptstyle d}}{{\sim}}\text{density}\ f(\cdot).$
If $\varepsilon_{n}$ has unit variance, $\sqrt{\varphi(\lambda_{n})}$ is in fact the conditional volatility of $y_{n}$. Here, the function $\varphi:\mathbb{R}\to\mathbb{R}^{+}$ is monotonic and differentiable, referred to as the link function. Moreover, if $\varphi$ is the identity mapping (restricted to $\mathbb{R}^{+}$), then the time-varying parameter $\lambda_{n}$ directly characterizes the conditional variance of $y_{n}$. According to Buccheri et al., (2021), the score in this case can be expressed as
$\nabla_{n}=\frac{-\varphi^{\prime}(\lambda_{n})}{2\varphi(\lambda_{n})}\left[1+\frac{f^{\prime}\left(\frac{y_{n}}{\sqrt{\varphi(\lambda_{n})}}\right)}{f\left(\frac{y_{n}}{\sqrt{\varphi(\lambda_{n})}}\right)}\frac{y_{n}}{\sqrt{\varphi(\lambda_{n})}}\right].$
In this section, we set $\varphi(x)={\rm e}^{2x}$, which implies that $\lambda_{n}:=\ln\sigma_{n}$ is the log-volatility. We assume that $f$ is the density of Generalized Error Distribution (GED) with parameter $\nu$, that is
$f(x)=\frac{1}{2^{1+1/\nu}\Gamma(1+1/\nu)}\exp\left(-\frac{|x|^{\nu}}{2}\right).$
In this case, the score is given by
$\nabla_{n}=\frac{\nu}{2}\left|\frac{y_{n}}{{\rm e}^{\lambda_{n}}}\right|^{\nu}-1=\frac{\nu}{2}|\varepsilon_{n}|^{\nu}-1.$
Since the Fisher information is a constant, we set $S(\lambda_{n})=1$. Thus, the dynamic of $\lambda_{n}$ is given by
$\lambda_{n}=\omega+\sum_{i=1}^{n}\phi_{i}\left[\lambda_{n-i}+\frac{\nu}{2}|\varepsilon_{n-i}|^{\nu}-1\right].$
Following the terminology in Harvey and Lange, (2017), we refer to this model as the Gamma-GED-EGARCH($\infty$). When $\nu=2$, the GED reduces to standard normal distribution, and this model corresponds to the EGARCH model in Nelson, (1991) but without asymmetry. Based on the previous definitions, we define the following sequence scaling process:
$\begin{split}&X_{t}^{(n)}=\frac{1}{\sqrt{n}}\sum_{i=1}^{\lfloor nt\rfloor}\left[\exp(\lambda_{i}^{(n)})\right]^{\frac{1-a_{n}}{\omega_{n}}\theta}\varepsilon_{i},\quad\varepsilon_{i}|\mathcal{F}_{i-1}\stackrel{{\scriptstyle d}}{{\sim}}\ \text{GED}(\nu),\\ &\Lambda_{t}^{(n)}:=\frac{1-a_{n}}{\omega_{n}}\theta\lambda^{(n)}_{\lfloor nt\rfloor},\quad\lambda_{t}^{(n)}=\omega_{n}+\sum_{i=1}^{t}\phi_{i}^{(n)}\left[\lambda_{t-i}^{(n)}+\frac{\nu}{2}|\varepsilon_{n-i}|^{\nu}-1\right],\end{split}$ (4.1)
where $\phi_{i}\sim K/i^{1+\alpha}$, $1-a_{n}\sim n^{-\alpha}$, $\omega_{n}\sim n^{-1/2}$.
Theorem 9.
For any $\alpha\in(1/2,1)$, as $n\to\infty$, the pair $(X_{t}^{(n)},\Lambda_{t}^{(n)})$ defined in (4.1) converges weakly to $(X_{t},\Lambda_{t})$ in the Skorohod space,
$\begin{split}&X_{t}=\int_{0}^{t}{\rm e}^{\Lambda_{s}}{\rm d}B_{s},\\ &\Lambda_{t}=\frac{1}{\Gamma(\alpha)}\int_{0}^{t}(t-s)^{\alpha-1}\kappa(\theta-\Lambda_{s}){\rm d}s+\frac{\sqrt{\nu}\kappa\mu}{\Gamma(\alpha)}\int_{0}^{t}(t-s)^{\alpha-1}{\rm d}W_{s}.\end{split}$ (4.2)
Where $B_{t},W_{t}$ are independent Brownian motions, $\kappa=\lim_{n\to\infty}\frac{\alpha n^{\alpha}(1-a_{n})}{K\Gamma(1-\alpha)}$, $\mu=\lim_{n\to\infty}\frac{\theta}{\sqrt{n}\omega_{n}}$.
Proof.
Denote $\mathbb{E}_{k-1}[\cdot]:=\mathbb{E}[\cdot|\mathcal{F}_{k-1}]$. By the property of GED, for any $p\geq 2$,
$\mathbb{E}_{k-1}[|\varepsilon_{k}|^{p}]=\frac{2^{p/\nu}}{\nu\Gamma(1+1/\nu)}\Gamma\left(\frac{p+1}{\nu}\right)<\infty.$
Thus,
$\mathbb{E}[|\nabla_{k}|^{p}]=\sum_{i=0}^{p}\binom{k}{i}(-1)^{k-i}\left(\nu/2\right)^{i}\mathbb{E}[|\varepsilon_{k}|^{\nu i}]<\infty.$
Especially, it can be easily computed that
$U^{2}(\lambda_{k})=\mathbb{E}_{k-1}[\nabla_{k}^{2}]=\frac{\nu^{2}}{4}\mathbb{E}_{k-1}[|\varepsilon|^{2\nu}]-\nu\mathbb{E}_{k-1}[|\varepsilon|^{\nu}]+1=\nu.$
The second equation then follows from Theorem 1. From the fact that $\varepsilon$ with its score are naturally uncorrelated, by Donsker’ invariance principle,
$\left(\sum_{i=1}^{\lfloor nt\rfloor}\frac{\varepsilon_{i}}{\sqrt{n}},\sum_{i=1}^{\lfloor nt\rfloor}\frac{\nabla_{i}}{\sqrt{n}}\right)\Rightarrow(B_{t},W_{t})$
are two independent Brownian motions. ∎
4.2 A extension as rough volatility approximations
The limit (4.2) is actually a rough volatility model, where $\Lambda_{t}$ characterizes the log-volatility of assets log-price $X_{t}$. Note that $\Lambda_{t}$ follows a rough OU process but is different with that in Gatheral et al., (2018) which characterizes log-volatility with a OU process driven by rough Brownian motion. Because in this volatility process (4.2), the kernel $(t-s)^{\alpha-1}$ not only emerge in noise but also in drift term. This setting is analogous with rough CIR in Jaisson and Rosenbaum, (2016) or rough Heston in El Euch and Rosenbaum, (2019).
However, before it can be considered a “useful” rough volatility model, there are still two important issues we need to deal with. First, the initial value of $(X_{t},\Lambda_{t})$ is zero in (4.2); second, the two Brownian motions are independent, which prevents the model from capturing the leverage effect in the market.
For the first one, we adopt the strategy in Wang and Cui, (2025), add a baseline in the dynamic of $\lambda_{t}^{(n)}$, that is
$\lambda_{t}^{(n)}=\omega_{t}^{(n)}+\sum_{i=1}^{t}\phi_{i}^{(n)}\left[\lambda_{t-i}^{(n)}+\frac{\nu}{2}|\varepsilon_{t-i}|^{\nu}-1\right],$
where
$\omega_{t}^{(n)}=\omega_{n}+\xi\omega_{n}\left(\frac{1}{1-a_{n}}\left(1-\sum_{s=1}^{t-1}\phi_{s}^{(n)}\right)-\sum_{s=1}^{t-1}\phi_{s}^{(n)}\right),$
with $\xi>0$. It result the limit $\Lambda_{t}$ becomes
$\Lambda_{t}=\theta\xi+\frac{1}{\Gamma(\alpha)}\int_{0}^{t}(t-s)^{\alpha-1}\kappa(\theta-\Lambda_{s}){\rm d}s+\frac{\sqrt{\nu}\kappa\mu}{\Gamma(\alpha)}\int_{0}^{t}(t-s)^{\alpha-1}{\rm d}W_{s}.$
[…]
$g(x)=\frac{(\rho+\zeta)^{2}}{2}|x|{\rm e}^{-\rho x-\zeta|x|}.$
In the framework of qausi-score driven model, it recovers the asymmetric structure analogous with the EGARCH model, that is
$\lambda_{t}^{(n)}=\omega_{n}+\sum_{i=1}^{t}\phi_{i}^{(n)}\left[\lambda_{t-i}^{(n)}+\rho\varepsilon_{t-i}+\zeta\left(|\varepsilon_{t-i}|-\sqrt{2/\pi}\right)\right].$
In summary, we consider the following sequence scaling qausi-score driven long memory volatility model,
$\begin{split}&X_{t}^{(n)}=X_{0}-\frac{1}{2n}\sum_{i=1}^{\lfloor nt\rfloor}\left[\exp(2\lambda_{i}^{(n)})\right]^{\frac{1-a_{n}}{\omega_{n}}\theta}+\frac{1}{\sqrt{n}}\sum_{i=1}^{\lfloor nt\rfloor}\left[\exp(\lambda_{i}^{(n)})\right]^{\frac{1-a_{n}}{\omega_{n}}\theta}\varepsilon_{i},\quad\varepsilon_{i}|\mathcal{F}_{i-1}\stackrel{{\scriptstyle d}}{{\sim}}\ \mathcal{N}(0,1),\\ &\Lambda_{t}^{(n)}:=\frac{1-a_{n}}{\omega_{n}}\theta\lambda^{(n)}_{\lfloor nt\rfloor},\quad\lambda_{t}^{(n)}=\omega_{t}^{(n)}+\sum_{i=1}^{t}\phi_{i}^{(n)}\left[\lambda_{t-i}^{(n)}+\rho\varepsilon_{t-i}+\zeta\left(|\varepsilon_{t-i}|-\sqrt{2/\pi}\right)\right].\end{split}$ (4.3)
It can be easily verified that the limit of (4.3) is the following rough volatility model:
$\begin{split}&X_{t}=X_{0}-\frac{1}{2}\int_{0}^{t}{\rm e}^{2\Lambda_{s}}{\rm d}s+\int_{0}^{t}{\rm e}^{\Lambda_{s}}{\rm d}B_{s},\\ &\Lambda_{t}=\theta\xi+\frac{1}{\Gamma(\alpha)}\int_{0}^{t}(t-s)^{\alpha-1}\kappa(\theta-\Lambda_{s}){\rm d}s+\frac{\sqrt{\rho^{2}+\zeta^{2}}\kappa\mu}{\Gamma(\alpha)}\int_{0}^{t}(t-s)^{\alpha-1}{\rm d}W_{s},\\ &\text{Cov}(B_{t},W_{t})=\rho t.\end{split}$ (4.4)
4.3 Numerical simulation
Based on the convergence result, we will simulate the time series (4.3) to approximate rough volatility model (4.4). Accordingly, the simulated paths can be employed to price a variety of options via Monte Carlo methods. In this part, the parameter choices are aligned with those adopted in prior numerical studies, such as Callegaro et al., (2021), Ma and Wu, (2022) and Wang and Cui, (2025), these values are:
$\begin{split}&X_{0}=\log 100,\quad\Lambda_{0}=\log\sqrt{0.0392},\quad\rho=-0.681,\\ &\kappa=0.1,\quad\theta=\log\sqrt{0.3156},\quad\sqrt{\rho^{2}+\zeta^{2}}\mu=0.331,\quad\alpha=0.62.\end{split}$
Note that the value of $\Lambda_{0}$ and $\theta$ implies that volatility is begin with $\sqrt{0.0392}\approx 19.80\%$, and the long trem level is $\sqrt{0.3156}\approx 56.18\%$. To visualize the volatility behavior, we first simulate Equation (4.3) over a time horizon $T=5$ with $n=1000$, varying the roughness parameter $\alpha$. Figure 1 shows that the path of the volatility $\sigma_{t}=\exp(\Lambda_{t})$ by simulating (4.3) for different $\alpha$.
Figure 1: Volatility paths for different $\alpha$ values
It can be observed that as $\alpha$ increases, the volatility paths exhibit greater smoothness, and the mean-reverting behavior becomes more evident. In other words, a smaller $\alpha$ not only implies rougher paths but also corresponds to a slower mean-reversion, which is consistent with the observations reported in Gatheral et al., (2018).
We proceed to price a range of path-independent and path-dependent options, including European, Asian, Lookback, and Barrier options, using Monte Carlo simulations. The Algorithm 1 outlines the computational steps for simulating asset paths and pricing these options.
Algorithm 1 Monte Carlo Option Pricing with EGARCH($\infty$) Approximation
Input: Model parameters ($\alpha$, $\kappa$, $\theta$, $\mu$, $\rho$), initial value ($X_{0}$, $\Lambda_{0}$), number of simulated paths $M$, time steps per year $n$, time horizon $T$, strike price $K$, Up-In barrier $B_{u}$, Down-Out barrier $B_{d}$.
Output: Option prices and standard errors.
Initialization: Compute discrete model parameters:
$\phi_{1}=1-\frac{1}{\Gamma(1-\alpha)\cdot 2^{\alpha}},\quad\phi_{i}=\frac{1}{\Gamma(1-\alpha)}\left(\frac{1}{i^{\alpha}}-\frac{1}{(i+1)^{\alpha}}\right)\text{ for }i\geq 2.$
$\omega_{n}=\frac{\theta}{\mu\sqrt{n}}$, $a_{n}=1-\kappa n^{-\alpha}$, $\xi=\frac{\Lambda_{0}}{\theta}$.
Key step: Generate $M$ asset price paths $S_{t}={\rm e}^{X_{t}}$ by simulating Equation (4.3).
for each path do
  Compute various option payoffs $V_{i}$:
  • European: $\max(S_{T}-K,0)$ for call, $\max(K-S_{T},0)$ for put
  • Asian: $\max(\frac{1}{nT}\sum_{i=1}^{nT}S_{t_{i}}-K,0)$ for call, $\max(K-\frac{1}{nT}\sum_{i=1}^{nT}S_{t_{i}},0)$ for put
  • Lookback: $\max(\max_{i=0,\ldots,nT}S_{t_{i}}-K,0)$ for call, $\max(K-\min_{i=0,\ldots,nT}S_{t_{i}},0)$ for put
  • Barrier: $\max(S_{T}-K,0)1_{\max S_{t_{i}}\geq B_{u}}$ for call, $\max(K-S_{T},0)1_{\min S_{t_{i}}>B_{d}}$ for put
end for
return option price estimates $V=\frac{1}{M}\sum_{i=1}^{M}V_{i}$, and standard errors $\sqrt{\frac{\sum_{i=1}^{M}(V_{i}-V)^{2}}{(M-1)M}}$.
Table 1 reports the option prices computed via Algorithm 1, with standard errors shown in parentheses. We set $M=10^{5}$, $n=500$, $T=1$, with strike prices $K=80,90,100,110,120$. For the barrier options, the barrier level is set to $B_{u}=110$ for the Up-In call and $B_{d}=90$ for the Down-Out put.
Table 1: Option pricing results under the rough volatility model (4.4)
Strike European Asian Lookback Barrier
Call Put Call Put Call Put Call Put
80 21.4331 1.4731 20.0771 0.1216 37.2152 2.6954 18.5647 0.0000
(0.0613) (0.0135) (0.0371) (0.0027) (0.0477) (0.0173) (0.0667) (0.0000)
90 13.9924 4.0324 11.1154 1.1600 27.2152 7.3847 13.0543 0.0000
(0.0537) (0.0237) (0.0328) (0.0097) (0.0477) (0.0280) (0.0552) (0.0000)
100 8.4271 8.4671 4.7219 4.7665 17.2152 15.4721 8.2995 0.1694
(0.0439) (0.0350) (0.0237) (0.0205) (0.0477) (0.0341) (0.0440) (0.0030)
110 4.7040 14.7440 1.5043 11.5489 9.4057 25.4721 4.7040 1.0715
(0.0336) (0.0454) (0.0137) (0.0299) (0.0419) (0.0341) (0.0336) (0.0103)
120 2.4492 22.4892 0.3638 20.4084 4.8067 35.4721 2.4492 2.8810
(0.0244) (0.0537) (0.0066) (0.0352) (0.0321) (0.0341) (0.0244) (0.0204)
To accelerate the path generation process, we employ two techniques: parallel computing and FFT-based convolution. Parallel computing distributes the simulation of independent paths across multiple processors, achieving near-linear speedup. The FFT-based approach reformulates the convolution operation in Equation (4.3) as:
$\sum_{i=1}^{t}\phi_{i}^{(n)}\left[\lambda_{t-i}^{(n)}+\rho\varepsilon_{t-i}+\zeta\left(|\varepsilon_{t-i}|-\sqrt{2/\pi}\right)\right]=\left(\Phi^{(n)}*\Psi^{(n)}\right)_{t},$
where $\Phi^{(n)}:=(\phi_{i}^{(n)})_{i=1,\ldots,nT}$ and $\Psi^{(n)}:=\left(\lambda_{i}^{(n)}+\rho\varepsilon_{i}+\zeta\left(|\varepsilon_{i}|-\sqrt{2/\pi}\right)\right)_{i=0,\ldots,nT-1}$. By applying Fast Fourier Transform (FFT), the computational complexity is reduced from $\mathcal{O}((nT)^{2})$ to $\mathcal{O}(nT\log nT)$, significantly improving efficiency for large $n$ and $T$. Table 2 compares the average computational time for generating a single path with and without FFT acceleration, demonstrating substantial performance gains as $n$ increases.
Table 2: Computation time with and without FFT acceleration (in seconds)
n(T=1)n\ (T=1) 100 250 500 1000 2500 5000
With FFT 0.0029 0.0073 0.0163 0.0383 0.2952 1.0789
Without FFT 0.0104 0.0586 0.2333 0.9208 5.8149 23.1063
5 Conclusion
This study reveals the connection between long memory score-driven models and the rough OU process, the former offering statistical robustness and the latter exhibiting rich dynamic behavior, especially in volatility. Specifically, we find that when $\alpha\in(1/2,1)$, the coefficients of the lag terms decay at a power-law rate of order $1+\alpha$, and the limiting process is driven by a fractional Brownian motion with Hurst parameter $H=\alpha-1/2$, which implies roughness. This result is inspired by the work of Jaisson and Rosenbaum, (2016) on scaling limits of Hawkes processes. In our proof, however, we introduce a novel approach based on GFOs. Leveraging their continuity in Hölder spaces, we directly establish the convergence of the parameter process in the corresponding Hölder space, leading to a more direct argument.
We apply our theoretical results to volatility modeling. When the time-varying parameter governed by the score-driven mechanism represents the log-volatility, the model converges to a new rough volatility model in which the log-volatility follows a rough OU process. More precisely, it is a process obtained by applying a Riemann-Liouville fractional operator of order $\alpha$ to an OU process, which differs from that in Gatheral et al., (2018).
In the numerical experiments, we approximate this rough volatility model using an EGARCH($\infty$) specification. As a quasi score-driven model, it incorporates the score of asymmetric distributions can thus capture the leverage effect. In this sense, our results may be seen as a long memory extension of Nelson, (1990)’s result for AR(1)(1) Exponential ARCH. The proposed Monte Carlo algorithm, optimized through FFT-based convolution and parallel computing, achieves significant computational efficiency, making it suitable for pricing diverse options, including European, Asian, Lookback, and Barrier options.
The results of this paper can be further developed in future work. In fact, the derivation of the OU-type limiting process relies on the assumption that the second-order conditional moment of the score, denoted by $U$, is asymptotically constant. While this condition is not difficult to satisfy—for instance, it always holds when the scaling function $S$ is chosen as the 1/2-1/2 power of the Fisher information—we believe that the results can be extended to more general forms of $U$, yielding corresponding limits of the form (2.8). Such an extension would require new techniques, as the GFO approach used here would break down: the limit process in this case cannot be represented as a GFO applied to the associated diffusion.
Acknowledgements
The authors express gratitude to Yingli Wang for insightful discussions and valuable suggestions, which significantly improved the quality of this work.